# ACCESS: Assurance Case Centric Engineering of Safety-critical Systems

### v1

@@ 5 Case Study
In this section, we evaluate the ACCESS methodology by applying it to a development process of an Autonomous Underwater Vehicle (AUV). The assurance case for the AUV is developed based on an integrated RoboChart miyazawa2016robochart model, which includes the architecture of the AUV and the behaviour of its controllers, discussed in foster2020formal . We also show that it is possible, with the help of ACME, to automate the evaluation of model based assurance cases and have the assurance case drive the development of the system.
@@ 5.2 ACCESS Step 2
The overall architecture of the AUV is modelled (task S3) using the RoboChart language miyazawa2016robochart , shown in Figure 11. The robotic platform (AUV_Platform) acts as an abstraction layer for the hardware, and provides shared variables for sensors, actuators and events. The operator, which can be a human or navigation system, provides instructions to the LRE (LRE_Ctrl - Last Response Engine) to support execution of tasks, such as requesting a particular heading and velocity. The LRE sits between the operator (AUV_Operator) and the autopilot component (AUV_Autopilot). The LRE’s job is to avoid hazardous behaviours, such as getting too close to an obstacle, or entering Object Proximity Exclusion Zones (OPEZs), and engaging evasive manoeuvres if necessary. The autopilot controls the AUV actuators, and takes advice only from the LRE.
@@ 5.3 ACCESS Step 3
To validate the FMEDA, we aim to achieve the target SPFM (assume we aim to achieve 90%). In order to do this automatically in ACME, in the assurance case, we create an ArtifactPackage (discussed in Section 2.3), and in it create an Artifact, and we refer to the Excel spreadsheet from the Artifact, which can be later used as evidence to our safety argument in the assurance case. As shown in Figure 13, in the references section, we configure the “Model Type” (as Excel spreadsheet), the “Docuemtn” (to the location of the Excel spreadsheet), and the “Metadata” (empty in this example). Within ACME, we use ImplementationConstraints to store model validation rules in Artifact elements so that such rules can be executed when we evaluate the assurance case. In Figure 13, in the “Constraint” section, we use the rule in Listing 1 (written in the Epsilon Object Language – EOL kolovos2006epsilon ) to check if the FMEDA fulfils our requirement for the target SPFM value. During the development process, when the FMEDA changes, ACME can automatically detect this change by the execution of the query and compute the SPFM automatically, then show the users if the hardware design fails to meet the target SPFM values.
@@ 5.4.2 Behaviour model for the LRE
We now model the behaviour of the LRE (task S5). We create a state machine for the LRE, shown in Figure 15. It implements the LRE’s behavioural requirements and specifies the conditions on switching to different operation modes. The following definitions and functions appear in the state machine miyazawa2016robochart : vel (velocity of the AUV), inOPEZ (if the AUV is in an OPEZ), CDA (Closest Distance of Approach), StaticObsHorizDist and StaticObsVertDist (shortest distance allowed to an obstacle horizontally and vertically), MinSafeDist (minimal overall safe distance), cdyn (closest dynamic obstacle), cstc (closest static obstacle), hdist() (horizontal distance to an obstacle), vdist() (vertical distance to an obstacle), odist() (overall distance to an obstacle).
@@
[…]
@@ 5.5.1 Trace to EMF models
We then attach the model validation rule in Listing 2 in the “Constraint” section. In this rule we check that there are at least 3 transitions from MOM, named “t4”, “t5” and “t6” (shown in Figure 15), which are triggered when there are potential collision risks. For readability we only show the queries for Transition “t4”. The user can evaluate the query inside the dialog by pressing the “Query” button. ACME will load the model specified in the “Reference” section and execute the query, the result of which is displayed in the “Query Result” text field. It is to be noted that the validation rules do not have to be specified by EOL, in a separate publication wei2023automated , we illustrated the use of constraint natural language with model-based approach.
@@ 5.5.4 Transformation to Isabelle/SACM
Once a satisfactory assurance case is developed, a further step is to check the integrity of the overall assurance case formally using Isabelle/SACM. This is an optional step for ACCESS, but we would like to show that how formalism can be integrated with model-based assurance case to form a more convincing assurance case. To do this, we perform a model-to-text transformation to generate Isabelle/SACM notations from the assurance case automatically, which can be machine-check using the Isabelle server. In ACME, we use the Epsilon Generation Language (EGL) rose2008epsilon to implement the transformation, but the transformation can be generalised. Algorithm 1 shows the pseudo-code for generating Isabelle/SACM notations from a GSN module.
@@
22             if source.isTypeOf(GSN!Solution) then  output “Inference ” + element.name + “src ‹{@{ArtifactReference ” + source_name + “}}› tgt ‹{@{Claim” + target_name +"}}›" + “ ‹@{Claim ” + target_name + “ } is supported by @{ArtifactReference ” + source_name + “}.›” else
@@
A fragment of transformed Isabelle/SACM argument is shown in Figure 19, with some elements reordered to aid readability. Context LRE_C1 and Solution Sn3 both give rise to references to formal artifacts included from another Isabelle theory file foster2020formal , including the formalised LRE state machine. Several of the claims are left open, and so have the keyword needsSupport attached, which can be checked to ensure that all branches of the argument are developed. An Inference, I1, connects the formalised Sn3 (the source) to the Claim C7_c (the target). The remainder of the claims, inferences, and solutions are similarly mapped.
@@ 5.7 ACCESS Step 7
  • Consequently, the safety analysis for the sensors can be updated accordingly, which we discussed in a recently published work wei2022designing ; wei2023decisive .
@@ 6.1 RQ1: Efficiency
The experiment procedure is as follows. We asked Participant A to engineer Subsystem A and Subsystem B by following ACCESS with a complete manual approach with no model-based tool support (i.e. all artifacts produced are text-based documents). We then asked Participant B to engineer Subsystem A and Subsystem B (by following ACCESS) with an model-based approach but without the support from ACME, Participant B chose to use EMF steinberg2008emf and Robochart Miyazawa2019-RoboChart as the modeling platform and chose Eclipse Epsilon kolovos2008epsilon as the model management framework. We then asked Participant C to engineer Subsystem A and Subsystem B (by following ACCESS) with the tool support from ACME. We recorded the effort it took for all three participants to engineer Subsystem A and Subsystem B in terms of time. The experiment results are shown in Table 2.
@@ 6.2 RQ2: Generality
In addition, throughout the development process using ACCESS with model-based support, we found that it was often necessary to “link” models defined using different modeling technologies. In our work, this is achieved by exploiting the facilities provided by the Structured Assurance Case Metamodel (SACM). However, other model-based approaches, for example, the use of a weaving model hawkins2015weaving may also be adopted to realise the links among heterogeneous models.
In addition, we would like to point out that, some components of System B were developed in conformance to ISO 26262 as SEooCs (Safety Element out of Context) iso26262 following the development process defined in ISO 26262. From our experience the ACCCESS methodology integrated seamlessly with the development of SEooCs which would be certified against ISO 26262.
@@ 6.4 RQ4: Scalability
Our last evaluation is on the scalability of ACME, although ACME is the secondary contribution of our work, we report our findings in the scalability of the tools. Our evaluation was performed on the premise that the majority of the models used in our development process are EMF models, with a mixture of Simulink models, Excel spreadsheets, formal models and JSON models, which are identified in Section 6.2. To evaluate the scalability of ACME, we selected 5 data sets as shown in Table 3. It is to be noted that the in our end result systems, the maximum number of model elements we have in our collection of models was 5689 (Set3). We made duplicates of our models and put them together to form Set4 and Set5 to evaluate the scalability of ACME in different order of magnitudes. We found that ACME suffered from scalability issues from Set4 and would not load Set5 due to memory overflow. This is typically caused by the fact that ACME need to load EMF models in their entirety before any queries can be performed on them, which is an existing issue discovered in various studies barmpis2013hawk ; wei2016partial ; shah2014framework .
@@ 7 Related Work
There are a number of assurance case works and tools that promote automation by adopting MDE, such as AdvoCATE denney2017tool , D-Case Editor matsuno2010dependability , ASCE netkachova2014tool , Astah GSN larrucea2017supporting , and CertWare barry2011certware .
@@
NASA has developed a powerful graphical tool AdvoCATE denney2017tool based on Eclipse EMF for AC generation, management, and evaluation. MDE is applied in two aspects; GSN is extended with a formal syntax to support the syntactic checks; and the evidence is generated by exploiting FM; it also uses a formal foundation for a lightweight semantic checking based on the metadata attached to the nodes though it is not to encode the whole argument in a formal machine-checkable language. The work follows the process of pattern design, instantiation, and formal verification. The AC pattern designed is generic without specific application constraints to the systems, and is split into two levels, from hazard mitigation to safety requirements, and from safety requirements down to the evidence. The claims at the second level, i.e. the safety requirements, are formalized manually, and verified by invoking the AUTOCERT tool Denney2008 . The instantiation requires engineers to identify the logic relationship among system data and between AC nodes and system data. The mapping is represented as a table to facilitate the automatic instantiation.
@@
Hawkins et al. proposed a model-based approach for generating Assurance Cases hawkins2015weaving . Model-based GSN assurance case patterns are used as a basis. The work exploits the concept of weaving models del2006weaving that represent the links between metamodels. With the weaving models, the approach allows the instantiation of GSN patterns, which automatically instantiate weaves (by means of model transformation) system information into an assurance case based on the links in the weaving models. The links between assurance case elements and system data can be updated automatically when the system design changes because the links are built between the metamodels instead of specific system data. However, apart from injecting system information into an assurance case, no traceability support from an assurance case to engineering artifacts is provided. Consequently, the validation and verification of assurance case is not covered.
@@
lin2016framework is also based on a pre-defined pattern and pre-organised data. The software development process is modelled. The process activities’ results are software artifacts whose relationships are shown through the input and output of processes. For example,the contribution of software to the hazard derives the software safety requirements. Thus, the relationships of different classes of data can be extracted from the development process models. The pattern is designed specifically for software safety with consideration of software contribution to the hazard. The work adds the syntax of GSN to allow automatic instantiation of the software artefacts. With the help of a relationship model, the impact of artefact change, e.g. the change of the software contribution to a hazard, can be identified automatically in a AC structure for a convenient review. However, since the software is not required to be developed with MDE, the automatic verification of AC is not considered in this work.
@@
Utsunomiya et al. utsunomiya2018tool developed a tool for constructing assurance cases by reading architecture models, quality properties, and risk measurements written in Extensible Markup Language (XML). Assurance cases are generated which conform to SACM v1.0 XML Metadata Interchange schema definition. The tool does not cover the evidence generation of assurance cases. The effectiveness of the method has been insufficiently evaluated to show whether it can be generalised.
@@
Prokhorova et al. prokhorova2015facilitating provide a solution for formalizing sub-claims and verifying with formal method tools. The system properties represented by sub-claims are categorised into eight classes such as temporal properties, timing properties, etc. Patterns are established for each class with different formal method verification solutions. The combination of verification tools shows the necessity of multiple formal methods, also referred to as integrated formal methods paige1997meta .
@@
Gleirscher et al. gleirscher2019evolution proposed to model assurance cases formally for autonomous robot systems following the pattern and instantiation process. The purpose is to cope with the AC evolution during the system development instead of the operation. Two assurance case patterns are designed, that cover the construction pattern for the system specification phase, and the extension pattern for the system implementation/ refinement phase. Both are instantiated with the system models. Since the system is required to be modelled in a formal language, there is no need for a process of formalisation of system models, and formal method is a natural choice for assurance case verification. However, the two-phase patterns are specific to the assumption/guarantee (A/G)-style reasoning, therefore not suitable for other RAS systems that do not follow A/G-style reasoning.
@@
Calinescu et al. calinescu2017engineering proposes the dynamic verification of self-adaptive systems at runtime, but no tool support has been provided. The work proposed the assurance case pattern for self-adaptive systems which is instantiated with formalised system requirements, and further verified by model checkers, which facilitates the automatic co-evolution of system and assurance case by avoiding the manual process of system model formalization.
@@
Gacek et al. gacek2014resolute proposed a new domain-specific language, Resolute, which is also a tool for building assurance cases based on AADLas55062004architecture models. The generation of assurance cases consists of two steps. Firstly, the top-level claim is defined formally by the engineers in first-order predicates where AADL models are queried. Secondly, the engineers decompose the top-level claim quantitatively. An example could be that the sum of message delay of threads is bounded while the top-level claim is the message delay of the process is bounded. Then the AADL model will be queried through the assurance case and checked automatically by model checking. The claims and rules for two steps are recorded in the Resolute library for reuse. The way to integrate Resolute assurance case with AADL model is to add a “prove" statement within AADL models. As the assurance case is written in the same developing environment of AADL models, system model to support assurance case claims will be queried by Resolute models, not transformed or instantiated. So the Resolute assurance case is directly integrated into the system architecture model. This integration enables the traceability and consistency between the system model and its assurance cases, and facilitates the automatic co-evolution of system and assurance case.But this integration on the other hand limits the solution application to other system modelling languages. Since ACME is based on the EMF and the Epsilon framework, it is possible to support various modelling languages (which is demonstrated in Section 5). It is noted that Resolute can only be applied to the system architecture level, rather than the implementation, otherwise it will incur state explosion. Also, the claim of the use case is security properties, where the architectural data are sufficient as assurance case inputs. However, it will require more data (e.g. hazard models, safety-related function models) for safety property argumentation; and Resolute may not be capable of describing those properties properly.
@@
Rushby conceived of an evidential tool bus Rushby2005-ETB that would allow integration of various verification tools to provide evidence to an assurance case, an idea that was later realised by Cruanes et al. Cruanes2013 . Isabelle is also an evidential tool bus, and its connection to ACME allows linking to formal evidence.

### v2

@@ 5 Case Study
In this section, we evaluate the ACCESS methodology by applying it to a development process of an Autonomous Underwater Vehicle (AUV). The assurance case for the AUV is developed based on an integrated RoboChart [42] model, which includes the architecture of the AUV and the behaviour of its controllers, discussed in [20]. We also show that it is possible, with the help of ACME, to automate the evaluation of model based assurance cases and have the assurance case drive the development of the system.
@@ 5.2 ACCESS Step 2
The overall architecture of the AUV is modelled (task S3) using the RoboChart language [42], shown in Figure 11. The robotic platform (AUV_Platform) acts as an abstraction layer for the hardware, and provides shared variables for sensors, actuators and events. The operator, which can be a human or navigation system, provides instructions to the LRE (LRE_Ctrl - Last Response Engine) to support execution of tasks, such as requesting a particular heading and velocity. The LRE sits between the operator (AUV_Operator) and the autopilot component (AUV_Autopilot). The LRE’s job is to avoid hazardous behaviours, such as getting too close to an obstacle, or entering Object Proximity Exclusion Zones (OPEZs), and engaging evasive manoeuvres if necessary. The autopilot controls the AUV actuators, and takes advice only from the LRE.
@@ 5.3 ACCESS Step 3
To validate the FMEDA, we aim to achieve the target SPFM (assume we aim to achieve 90%). In order to do this automatically in ACME, in the assurance case, we create an ArtifactPackage (discussed in Section 2.3), and in it create an Artifact, and we refer to the Excel spreadsheet from the Artifact, which can be later used as evidence to our safety argument in the assurance case. As shown in Figure 13, in the references section, we configure the “Model Type” (as Excel spreadsheet), the “Docuemtn” (to the location of the Excel spreadsheet), and the “Metadata” (empty in this example). Within ACME, we use ImplementationConstraints to store model validation rules in Artifact elements so that such rules can be executed when we evaluate the assurance case. In Figure 13, in the “Constraint” section, we use the rule in Listing 1 (written in the Epsilon Object Language – EOL [33]) to check if the FMEDA fulfils our requirement for the target SPFM value. During the development process, when the FMEDA changes, ACME can automatically detect this change by the execution of the query and compute the SPFM automatically, then show the users if the hardware design fails to meet the target SPFM values.
@@ 5.4.2 Behaviour model for the LRE
We now model the behaviour of the LRE (task S5). We create a state machine for the LRE, shown in Figure 15. It implements the LRE’s behavioural requirements and specifies the conditions on switching to different operation modes. The following definitions and functions appear in the state machine [42]: vel (velocity of the AUV), inOPEZ (if the AUV is in an OPEZ), CDA (Closest Distance of Approach), StaticObsHorizDist and StaticObsVertDist (shortest distance allowed to an obstacle horizontally and vertically), MinSafeDist (minimal overall safe distance), cdyn (closest dynamic obstacle), cstc (closest static obstacle), hdist() (horizontal distance to an obstacle), vdist() (vertical distance to an obstacle), odist() (overall distance to an obstacle).
@@
$reqMOM\left[\begin{array}[]{l}vel\leq 0.1\land odist(cdyn)>7.5\\ \land odist(cstc)>0.3\land\neg\textit{inOPEZ}\end{array}\right]$
@@ 5.5.1 Trace to EMF models
We then attach the model validation rule in Listing 2 in the “Constraint” section. In this rule we check that there are at least 3 transitions from MOM, named “t4”, “t5” and “t6” (shown in Figure 15), which are triggered when there are potential collision risks. For readability we only show the queries for Transition “t4”. The user can evaluate the query inside the dialog by pressing the “Query” button. ACME will load the model specified in the “Reference” section and execute the query, the result of which is displayed in the “Query Result” text field. It is to be noted that the validation rules do not have to be specified by EOL, in a separate publication [64], we illustrated the use of constraint natural language with model-based approach.
@@ 5.5.4 Transformation to Isabelle/SACM
Once a satisfactory assurance case is developed, a further step is to check the integrity of the overall assurance case formally using Isabelle/SACM. This is an optional step for ACCESS, but we would like to show that how formalism can be integrated with model-based assurance case to form a more convincing assurance case. To do this, we perform a model-to-text transformation to generate Isabelle/SACM notations from the assurance case automatically, which can be machine-check using the Isabelle server. In ACME, we use the Epsilon Generation Language (EGL) [51] to implement the transformation, but the transformation can be generalised. Algorithm 1 shows the pseudo-code for generating Isabelle/SACM notations from a GSN module.
@@
22             if source.isTypeOf(GSN!Solution) then  output “Inference ” + element.name + “src ‹{@{ArtifactReference ” + source_name + “}}› tgt ‹{@{Claim” + target_name +"}}›" + “ ‹@{Claim ” + target_name + “ } is supported by @{ArtifactReference ” + source_name + “}.›”
23             else
@@
A fragment of transformed Isabelle/SACM argument is shown in Figure 19, with some elements reordered to aid readability. Context LRE_C1 and Solution Sn3 both give rise to references to formal artifacts included from another Isabelle theory file [20], including the formalised LRE state machine. Several of the claims are left open, and so have the keyword needsSupport attached, which can be checked to ensure that all branches of the argument are developed. An Inference, I1, connects the formalised Sn3 (the source) to the Claim C7_c (the target). The remainder of the claims, inferences, and solutions are similarly mapped.
@@ 5.7 ACCESS Step 7
  • Consequently, the safety analysis for the sensors can be updated accordingly, which we discussed in a recently published work [62, 63].
@@ 6.1 RQ1: Efficiency
The experiment procedure is as follows. We asked Participant A to engineer Subsystem A and Subsystem B by following ACCESS with a complete manual approach with no model-based tool support (i.e. all artifacts produced are text-based documents). We then asked Participant B to engineer Subsystem A and Subsystem B (by following ACCESS) with an model-based approach but without the support from ACME, Participant B chose to use EMF [54] and Robochart [43] as the modeling platform and chose Eclipse Epsilon [34] as the model management framework. We then asked Participant C to engineer Subsystem A and Subsystem B (by following ACCESS) with the tool support from ACME. We recorded the effort it took for all three participants to engineer Subsystem A and Subsystem B in terms of time. The experiment results are shown in Table 2.
@@ 6.2 RQ2: Generality
In addition, throughout the development process using ACCESS with model-based support, we found that it was often necessary to “link” models defined using different modeling technologies. In our work, this is achieved by exploiting the facilities provided by the Structured Assurance Case Metamodel (SACM). However, other model-based approaches, for example, the use of a weaving model [26] may also be adopted to realise the links among heterogeneous models.
In addition, we would like to point out that, some components of System B were developed in conformance to ISO 26262 as SEooCs (Safety Element out of Context) [28] following the development process defined in ISO 26262. From our experience the ACCCESS methodology integrated seamlessly with the development of SEooCs which would be certified against ISO 26262.
@@ 6.4 RQ4: Scalability
Our last evaluation is on the scalability of ACME, although ACME is the secondary contribution of our work, we report our findings in the scalability of the tools. Our evaluation was performed on the premise that the majority of the models used in our development process are EMF models, with a mixture of Simulink models, Excel spreadsheets, formal models and JSON models, which are identified in Section 6.2. To evaluate the scalability of ACME, we selected 5 data sets as shown in Table 3. It is to be noted that the in our end result systems, the maximum number of model elements we have in our collection of models was 5689 (Set3). We made duplicates of our models and put them together to form Set4 and Set5 to evaluate the scalability of ACME in different order of magnitudes. We found that ACME suffered from scalability issues from Set4 and would not load Set5 due to memory overflow. This is typically caused by the fact that ACME need to load EMF models in their entirety before any queries can be performed on them, which is an existing issue discovered in various studies [3, 66, 53].
@@ 7 Related Work
There are a number of assurance case works and tools that promote automation by adopting MDE, such as AdvoCATE [13], D-Case Editor [40], ASCE [45], Astah GSN [35], and CertWare [4].
@@
NASA has developed a powerful graphical tool AdvoCATE [13] based on Eclipse EMF for AC generation, management, and evaluation. MDE is applied in two aspects; GSN is extended with a formal syntax to support the syntactic checks; and the evidence is generated by exploiting FM; it also uses a formal foundation for a lightweight semantic checking based on the metadata attached to the nodes though it is not to encode the whole argument in a formal machine-checkable language. The work follows the process of pattern design, instantiation, and formal verification. The AC pattern designed is generic without specific application constraints to the systems, and is split into two levels, from hazard mitigation to safety requirements, and from safety requirements down to the evidence. The claims at the second level, i.e. the safety requirements, are formalized manually, and verified by invoking the AUTOCERT tool [15]. The instantiation requires engineers to identify the logic relationship among system data and between AC nodes and system data. The mapping is represented as a table to facilitate the automatic instantiation.
@@
Hawkins et al. proposed a model-based approach for generating Assurance Cases [26]. Model-based GSN assurance case patterns are used as a basis. The work exploits the concept of weaving models [12] that represent the links between metamodels. With the weaving models, the approach allows the instantiation of GSN patterns, which automatically instantiate weaves (by means of model transformation) system information into an assurance case based on the links in the weaving models. The links between assurance case elements and system data can be updated automatically when the system design changes because the links are built between the metamodels instead of specific system data. However, apart from injecting system information into an assurance case, no traceability support from an assurance case to engineering artifacts is provided. Consequently, the validation and verification of assurance case is not covered.
@@
[37] is also based on a pre-defined pattern and pre-organised data. The software development process is modelled. The process activities’ results are software artifacts whose relationships are shown through the input and output of processes. For example,the contribution of software to the hazard derives the software safety requirements. Thus, the relationships of different classes of data can be extracted from the development process models. The pattern is designed specifically for software safety with consideration of software contribution to the hazard. The work adds the syntax of GSN to allow automatic instantiation of the software artefacts. With the help of a relationship model, the impact of artefact change, e.g. the change of the software contribution to a hazard, can be identified automatically in a AC structure for a convenient review. However, since the software is not required to be developed with MDE, the automatic verification of AC is not considered in this work.
@@
Utsunomiya et al. [61] developed a tool for constructing assurance cases by reading architecture models, quality properties, and risk measurements written in Extensible Markup Language (XML). Assurance cases are generated which conform to SACM v1.0 XML Metadata Interchange schema definition. The tool does not cover the evidence generation of assurance cases. The effectiveness of the method has been insufficiently evaluated to show whether it can be generalised.
@@
Prokhorova et al. [50] provide a solution for formalizing sub-claims and verifying with formal method tools. The system properties represented by sub-claims are categorised into eight classes such as temporal properties, timing properties, etc. Patterns are established for each class with different formal method verification solutions. The combination of verification tools shows the necessity of multiple formal methods, also referred to as integrated formal methods [49].
@@
Gleirscher et al. [22] proposed to model assurance cases formally for autonomous robot systems following the pattern and instantiation process. The purpose is to cope with the AC evolution during the system development instead of the operation. Two assurance case patterns are designed, that cover the construction pattern for the system specification phase, and the extension pattern for the system implementation/ refinement phase. Both are instantiated with the system models. Since the system is required to be modelled in a formal language, there is no need for a process of formalisation of system models, and formal method is a natural choice for assurance case verification. However, the two-phase patterns are specific to the assumption/guarantee (A/G)-style reasoning, therefore not suitable for other RAS systems that do not follow A/G-style reasoning.
@@
Calinescu et al. [10] proposes the dynamic verification of self-adaptive systems at runtime, but no tool support has been provided. The work proposed the assurance case pattern for self-adaptive systems which is instantiated with formalised system requirements, and further verified by model checkers, which facilitates the automatic co-evolution of system and assurance case by avoiding the manual process of system model formalization.
@@
Gacek et al. [21] proposed a new domain-specific language, Resolute, which is also a tool for building assurance cases based on AADL[2] models. The generation of assurance cases consists of two steps. Firstly, the top-level claim is defined formally by the engineers in first-order predicates where AADL models are queried. Secondly, the engineers decompose the top-level claim quantitatively. An example could be that the sum of message delay of threads is bounded while the top-level claim is the message delay of the process is bounded. Then the AADL model will be queried through the assurance case and checked automatically by model checking. The claims and rules for two steps are recorded in the Resolute library for reuse. The way to integrate Resolute assurance case with AADL model is to add a “prove" statement within AADL models. As the assurance case is written in the same developing environment of AADL models, system model to support assurance case claims will be queried by Resolute models, not transformed or instantiated. So the Resolute assurance case is directly integrated into the system architecture model. This integration enables the traceability and consistency between the system model and its assurance cases, and facilitates the automatic co-evolution of system and assurance case.But this integration on the other hand limits the solution application to other system modelling languages. Since ACME is based on the EMF and the Epsilon framework, it is possible to support various modelling languages (which is demonstrated in Section 5). It is noted that Resolute can only be applied to the system architecture level, rather than the implementation, otherwise it will incur state explosion. Also, the claim of the use case is security properties, where the architectural data are sufficient as assurance case inputs. However, it will require more data (e.g. hazard models, safety-related function models) for safety property argumentation; and Resolute may not be capable of describing those properties properly.
@@
Rushby conceived of an evidential tool bus [52] that would allow integration of various verification tools to provide evidence to an assurance case, an idea that was later realised by Cruanes et al. [11]. Isabelle is also an evidential tool bus, and its connection to ACME allows linking to formal evidence.